OpenStudy (anonymous):

(integral sign) 2/sqrt(4-x^2) dx using the substitution x= 2sin theta

OpenStudy (anonymous):

get \[u=2\sin(\theta)\] \[du=2\cos(\theta)d\theta\] and get \[\int\frac{\cos(\theta)}{\sqrt{4-4\sin^2(\theta)}}\]

OpenStudy (anonymous):

denominator is \[\sqrt{4(1-\sin^2(\theta))}=2\cos(\theta)\]

OpenStudy (anonymous):

since \[1-\sin^2(\theta)=\cos^2(\theta)\]

OpenStudy (anonymous):

making your integral \[\frac{1}{2}\int1d\theta\]

OpenStudy (anonymous):

pretty easy integral you get \[\frac{1}{2}\theta\]

OpenStudy (anonymous):

now since \[x=2\sin(\theta)\] you have \[\frac{x}{2}=\sin(\theta)\] and therefore \[\theta =\sin^{-1}(\frac{x}{2})\]

OpenStudy (anonymous):

so answer is \[\frac{1}{2}\sin^{-1}(\frac{x}{2})\] oh and don't forget plus C

OpenStudy (anonymous):

no i screwed up. i am off by the constant. it is \[2\sin^{-1}(\frac{x}{2})+C \]sorry